A population has mean 75 and standard deviation 12.

c) Random samples of size 121 are taken. Find the mean and standard deviation of

the sample mean.


1
Expert's answer
2021-09-07T03:05:38-0400

X ~ N ( µ = 75 , σ = 12 )

Part a)

Mean of the sample mean µ = µ = 75


Standard deviation of sample mean σ"="σ/"\sqrt{(n)}" "=" 12"\/" "\sqrt{121}"

"=" 1.0909

Partb;

Mean of the sample mean µ = µ = 75

Standard deviation of sample mean σ = σ"\/" "\sqrt{(n)}" "=" 12"\/" "\sqrt{400}"

=0.6